Big-O notation for eachpublic static double accumulate (double[] a) { double sum = 0.0; for (int i = 0; i < a.length; i++) sum += a[i]; return sum;
Answer #1

SOL:

1) The time complexity is O(a.length)

if we consider a.length =n

Then Big-O is O(n)

Explanation:

The Loop runs from O to a.length . so the no of iterations will be a.length

The time complexity is directly propotional to number of iterations

No of iterations will be a.length

if we consider a.length = n

So Big-O is O(n)

2) The time complexity is O(a.length)

if we consider a.length =n

Then Big-O is O(n)

Explanation:

The Loop runs from O to a.length .so the no of iterations will be a.length

The time complexity is directly propotional to number of iterations

No of iterations will be a.length

if we consider a.length = n

So Big-O is O(n)

3) The time complexity is O(n2) i.e Big-O is O(n2) if we consider a.length =n

Explanation:

The time complexity is directly propotional to number of iterations

For i=1 The no of iterations will be n-1 times

For i=2 The no of iterations will be n-2 times

for i=3 The no of iterations will be n-3 times

So the no of iterations will be

n-1+n-2+n-3+...+1

=n(n+1)/2

= n2+n/2

= O(n2)

4) The time complexity is O(a.length)

if we consider a.length =n

Then Big-O is O(n)

Explanation:

The Loop runs from O to a.length . so the no of iterations will be a.length

The time complexity is directly propotional to number of iterations

No of iterations will be a.length

if we consider a.length = n

So Big-O is O(n)